OpenStudy (anonymous):

Find the circumference. Leave your answer in terms of CIRCLE HAS A RADIUS OF 10

OpenStudy (anonymous):

Answer must be in terms of PI.

OpenStudy (anonymous):

i just left my answer in terms of

OpenStudy (saifoo.khan):

in terms of.. "blank"

OpenStudy (anonymous):

PI.

OpenStudy (anonymous):

the formula is : Circumference = \(\large 2\pi*radius=2\pi(10)=20\pi \)

OpenStudy (anonymous):

20PI

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

What about if the diameter is 4.1m? how would i solve that?

OpenStudy (anonymous):

the diameter is twice the radius... so take half of 4.1. that will be your radius...